# Low-resource Machine Translation: what for? who for? An observational study on a dedicated Tetun language translation service

### v1

@@ 4.2 Translation direction and text length
Using the source and target language codes present in each of the server logs, we extract summary information on the distribution by source and target language. We find that Tetun is more likely to be the target language (68% of requests), with English to Tetun translation most in demand (48% of requests), followed by Tetun to English (25% of requests), Portuguese to Tetun (11%), Indonesian to Tetun (9%), Tetun to Portuguese (4.6%), and Tetun to Indonesian (2.6%) – see Table 2.
@@
English to Tetun 48%
Portuguese to Tetun 11%
Indonesian to Tetun 9%
Total Tetun as Target 68%
Tetun to English 25%
Tetun to Portuguese 4.6%
Tetun to Indonesian 2.6%
Total Tetun as Source 32%
@@
Leveraging an open source Tetun tokenizer de Jesus and Nunes (2024a) for Tetun, and the NLTK library for English/Indonesian/Portuguese, we count the number of words in the input texts. We find that shorter text inputs are much more common, with a sharp decline in frequency as the text length increases. Overall, inputted text has a median of 8 words, with a median of 12 words when Tetun is the target of translation, and 5 words when Tetun is the source. A histogram of word count frequency is shown in Figure 4.
025507500 k10101010 k20202020 k30303030 kLength of input text (number of words)
@@ 4.3.1 Topic
Using the input text and its associated MT-translated result from the server logs, we estimate what topics are most or least in-demand for text translation. Our results for topic coverage are shown in Table 3. We include a comparison with topic coverage for the largest open Tetun dataset currently available, MADLAD-400 Kudugunta et al. (2023).
We find overall that "Science & research" is the topic most in demand, present in 34.2% of MT inputs when weighted by word count, followed by "Healthcare & medicine", "Education", and "Business & work & employment", all present in 20 to 25% of inputs. Overall, "Religion & spirituality" is the least frequent topic in MT inputs, which comes in contrast with the high representation of the religious domain in low-resource corpora Haddow et al. (2022).
We observe a large discrepancy between user-translated text and available corpora, where the main topics covered by available corpora (Government & socio-economic issues, 83.1%; Law & regulations, 28.2%) do not match with the main topics in demand by MT users.
@@
Science & research 34.17% 3.173.173.173.17%
Healthcare & medicine 23.90% 11.9011.9011.9011.90%
Education 22.82% 11.2811.2811.2811.28%
Business & work & empl.. 21.19% 6.186.186.186.18%
Daily life & personal exp. 17.1517.1517.1517.15% 8.668.668.668.66%
Government & socio. issues 16.1716.1716.1716.17% 83.09%
Law & regulations 8.908.908.908.90% 28.21%
Technology 8.068.068.068.06% 3.903.903.903.90%
Religion & spirituality 6.806.806.806.80% 14.5114.5114.5114.51%
Table 3: Topic distribution, where each document can cover several topics. Highlighted values indicate percentages above 20%.
0020202020404040406060606080808080100100100100TechnologyDaily Life & Personal ExperiencesHealthcare & MedicineLaw & RegulationsBusiness & Work & EmploymentGovernment & Socio-economic IssuesEducationScience & ResearchReligion & Spirituality3.93.93.93.98.668.668.668.6611.911.911.911.928.2128.2128.2128.216.186.186.186.1883.0983.0983.0983.0911.2811.2811.2811.283.173.173.173.1714.5114.5114.5114.518.068.068.068.0617.1517.1517.1517.1523.923.923.923.98.98.98.98.921.1921.1921.1921.1916.1716.1716.1716.1722.8222.8222.8222.8234.1734.1734.1734.176.86.86.86.8Percentage (%)MT textMADLAD
Figure 5: Comparison of topic coverage for text translated on the MT service with the largest openly available Tetun corpus (MADLAD-400 "clean" split)
@@ 4.3.2 Provenance
See Table 4. We observe a large dominance of "Education & research material", which alone represent more than half of MT inputs (55.9%). "Organisational & formal documents" come as a distant second at 18%, followed by "Conversation & correspondence" at 8.7%. This again stands in contrast with monolingual corpora, where over 61% of documents come from "News articles & press releases", while only 7.5% come from "Education & research material".
Overall, our findings on topic and provenance are consistent with the hypothesis that most users of this MT service are students (who are more likely than the general population to translate educational material related to science, education or healthcare) and with de Jesus and Nunes (2024c) who find that most Tetun text on the internet comes from the news domain, which is likely to overlap with the "Government & socio-economic issues" topic.
@@
Education & research material 55.90% 7.517.517.517.51%
Organisational & formal docs 18.0018.0018.0018.00% 17.2817.2817.2817.28%
News article & press release 10.2010.2010.2010.20% 61.16%
Conversation & correspond. 8.728.728.728.72% 4.754.754.754.75%
Literary fiction 4.264.264.264.26% 0.890.890.890.89%
Religious text 2.522.522.522.52% 8.158.158.158.15%
Other 0.400.400.400.40% 0.260.260.260.26%
Table 4: Provenance distribution in MT and MADLAD (highlighted values indicate percentages above 20%)
@@ 4.3.3 Genre
Overall, we observe that official (INL) spelling is only loosely followed by MT inputs, with a median of 71% of words spelled correctly, compared to a median of 82% in MADLAD. This is coherent with previous studies that note the high proportion of Tetun texts that do not follow the official spelling, including in the media Greksakova (2018). We also observe that variance in the rate of correct spelling is higher in MT inputs (standard deviation of 13%) than in MADLAD (standard deviation of 8%) – see Figure 6.
@@
4.4 Further comparison with monolingual corpora
Surprisal
To illustrate the lack of domain overlap between MT inputs and available corpora identified in Section 4.3, we leverage a pre-trained Tetun language model to compare surprisal on MT Tetun inputs with surprisal on monolingual corpora. We use the Goldfish Tetun model Chang et al. (2024), which is trained mostly (> 98%) on MADLAD Tetun data. We get an average surprisal of 6.1 on MT Tetun inputs, compared to an average of 4.3 for Labadain (where we exclude sentences that appear in MADLAD). This higher surprisal suggests that the MT inputs may contain vocabulary, structures, or contexts that are underrepresented in scraped monolingual corpora, potentially leading to lower quality translations or increased translation errors.
Comparing average surprisal per domain, we find that documents in the "Government & Socio-economic issues" and "Religion & spirituality" domains have a lower surprisal (5.74 and 5.86 respectively) than documents from the "Science & research" and Technology (6.42 and 6.69 respectively). This result can be explained by the domain mismatch between MT inputs and monolingual corpora uncovered in Section 4.3.
Semantic embeddings
Working with a sample of 1,000 translation logs and 1,000 MADLAD Tetun documents, we aim to get a semantic representation of documents for comparison. Given the lack of Tetun semantic embedding model, we work with the high-resource side (in English, Indonesian or Portuguese) for translation logs, and we translate the MADLAD Tetun data to English using MADLAD-3B. We use LASER encoders Team et al. (2022), which support English, Indonesian and Portuguese, to get a shared representation of both translation logs and translated MADLAD data. After embedding, we visualise our result using t-SNE Van der Maaten and Hinton (2008). We find overall a large degree of separation between translation logs and MADLAD monolingual Tetun corpora, shown in Figure 7.
Figure 7: t-SNE visual comparison of MT text vs MADLAD embeddings
@@ 5 Recommendations
Future NLP work. Our results seem to indicate that future work on Tetun MT should focus on: (1) broad domain coverage that would include texts about daily life, science and health, (2) accuracy when translating into Tetun rather than from Tetun, and (3) accuracy when translating very short documents (including individual words). This could start with an evaluation framework that finds a balance between these criteria. In particular, the sharp imbalance between the domain of available Tetun corpora and MT inputs leads us to recommend the systematic evaluation of MT accuracy in a wide variety of domains, and advise against measuring MT accuracy on domains that are over-represented in available corpora, as this would risk giving a poor representation of end-user experience.
Future Human-Computer Interaction (HCI) work. We recommend a focus on meeting the needs of students who use MT on their mobile device. Given that students seem to represent a large portion of MT users, and that learning outcomes are an important social good, future study of MT perception and improvement for students in low-resource contexts would be beneficial, both to the research community and end users. Similarly, given the high cost and low reliability of internet connections in Timor Foundation (2023), we recommend a heightened focus on MT inference on mobile devices, for example through further research into efficient MT, or through the creation of an open source framework for on-device MT that targets mobile devices [^6] Similar to the Bergamot / Browser MT project.
6 Limitations
Design decisions made by the volunteers who built this MT service might have impacted usage numbers. For example, apart from MT, the service includes a multilingual word dictionary (with entry definitions and example sentences), and a Tetun spell checker, which might skew the number of users upward, with potentially some users using the app or website without using the MT functionality. Similarly, the authors chose to only allow translation between Tetun and three higher-resource languages (English, Portuguese, Indonesian), which raises the question of how adding support for more languages would impact usage and balance between languages.
This paper focuses on Tetun, which might limit the applicability of our findings to other languages, given the variability of use cases, users needs and data availability within low-resource languages Liu et al. (2022). While some of our findings will be applicable to other low-resource languages, in particular those that are also lingua franca, they are less likely to be applicable to much lower-resourced languages, in particular endangered ones.
Lastly, we identify some potential shortcomings of our domain classification methodology: (1) some classes are overlapping (e.g. text about biology could fall under "Health & medicine" or "Science & technology"), which makes them harder to distinguish, (2) it does not identify the source of each text, e.g. a document about healthcare could be from a news article or from a research article, and (3) it works only on documents that longer than 5 words, but most translated documents are very short, thus not covered by our classification methodology.
@@ 7 Conclusion
We present an analysis of MT usage for Tetun, a low-resource language, based on server logs from a widely-used MT service. Our study reveals key insights into user demographics, device preferences, translation directions, and domain coverage. We find that a large share of users are likely students in Timor-Leste using mobile devices, with a preference for translating into Tetun rather than from it. Our analysis also uncovers a significant mismatch between the domains of user-translated text and existing Tetun corpora.
@@
Our recommendations include a stronger emphasis on domains that are not well covered by existing corpora (in particular domains that are important to education), accuracy for translation into Tetun, performance on very short texts, and the development of mobile-optimised MT solutions. For the HCI community, we suggest focusing on MT perception and improvement for students in low-resource contexts.
@@
This study provides a first observational analysis of real-world MT usage for a low-resource language. It contributes to a more nuanced understanding of the challenges and opportunities in this field. We hope these insights will guide future research and development efforts, ultimately leading to more effective and accessible MT solutions for speakers of low-resource languages.

### v4

@@ 4.2 Translation direction and text length
Using the source and target language codes present in each of the server logs, we extract summary information on the distribution by source and target language. We find that Tetun is more likely to be the target language (over 70% of requests), with English to Tetun translation most in demand (46.7% of requests), followed by Tetun to English (22.3% of requests), Portuguese to Tetun (12.4%), Indonesian to Tetun (11.4%), Tetun to Portuguese (4.5%), and Tetun to Indonesian (2.5%) – see Table 2. For all language pairs, the high-resource to low-resource direction is in higher demand than the low-resource to high-resource direction.
@@
English to Tetun 46.7 %
Portuguese to Tetun 12.4 %
Indonesian to Tetun 11.4 %
Total Tetun as Target 70.6 %
Tetun to English 22.3 %
Tetun to Portuguese 4.5 %
Tetun to Indonesian 2.5 %
Total Tetun as Source 29.4 %
@@
Leveraging an open source Tetun tokenizer de Jesus and Nunes (2024a) for Tetun, and the spaCy library [^8] https://spacy.io/, version 3.8.2, model xx_ent_wiki_sm for English/Indonesian/Portuguese, we count the number of words in the input texts. We find that shorter text inputs are much more common, with a sharp decline in frequency as the text length increases. This does not include hits to the tetun.org dictionary, as the dictionary works offline. Overall, inputted text has a median of 8 words, with a median of 12 words when Tetun is the target of translation, and 5 words when Tetun is the source. We note however that this distribution has a long tail, with over 12% of translations having 95 input words or more. A histogram of word count frequency is shown in Figure 4.
025507595+$0$ k5555 k10101010 k15151515 k20202020 k25252525 k30303030 k35353535 kLength of input text (number of words)
@@ 4.3 Domain coverage
In this section, we investigate what domains are most or least in demand by tetun.org users, and how they compare to domains present in available Tetun corpora, as represented by the MADLAD corpus. Given that domain similarity between training data and evaluation set is one of the main driver of MT accuracy Khiu et al. (2024), this investigation can inform future efforts on low-resource corpus gathering.
@@ 4.3.1 Topic
Using the input text and its associated MT-translated result from the server logs, we estimate what topics are most or least in-demand for text translation. Our results for topic coverage are shown in Table 3. We include a comparison with topic coverage for the largest monolingual Tetun dataset currently available, MADLAD-400 Kudugunta et al. (2023).
We find overall that “Science & research” is the topic most in demand, present in 34.2% of MT inputs when weighted by word count, followed by “Healthcare & medicine”, “Education”, and “Business & work & employment”, all present in 20 to 25% of inputs. Overall, “Religion & spirituality” is the least frequent topic in MT inputs, which contrasts with the high representation of the religious domain in most low-resource corpora Haddow et al. (2022).
We observe a large discrepancy between user-translated text and available corpora, where the main topics covered by MADLAD (Government & socio-economic issues, 83.1%; Law & regulations, 28.2%) do not match with the main topics in demand by MT users.
@@
Science & research 34.2% 3.23.23.23.2%
Healthcare & medicine 23.9% 11.9%
Education 22.8% 11.3%
Business & work & empl.. 21.2% 6.26.26.26.2%
Daily life & personal exp. 17.2% 8.78.78.78.7%
Government & socio. issues 16.2% 83.1%
Law & regulations 8.98.98.98.9% 28.2%
Technology 8.18.18.18.1% 3.93.93.93.9%
Religion & spirituality 6.86.86.86.8% 14.5%
Table 3: Topic distribution, where each document can cover several topics.
We further analyze topic coverage depending on whether the source text is in Tetun or whether Tetun is the target language. When Tetun is the target, topics such as “Science & research” (36.73%), “Healthcare & medicine” (25.52%), and “Education” (23.14%) dominate, reflecting user demand for translating high-resource content into Tetun, likely to support knowledge dissemination and accessibility. In contrast, when Tetun is the source language, translations are predominantly focused on “Daily life & personal experiences” (48.64%), indicating that users frequently translate everyday communication into high-resource languages. This highlights a divergence in translation use cases: translating into Tetun serves educational and professional purposes, whereas translating from Tetun often fulfills personal or practical communication needs.
@@ 4.3.2 Provenance
As illustrated in Table 4, we observe a large dominance of “Education & research material”, which alone represents more than half of MT inputs (55.9%). “Organizational & formal documents” comes as a distant second at 18%. This again stands in contrast with monolingual corpora, where over 61% of documents come from “News articles & press releases”, while only 7.5% come from “Education & research material”.
Our findings on topic and provenance (Tables  3 and 4) are consistent with the hypothesis that most tetun.org users are students (who are more likely than the general population to translate educational material related to science, education or healthcare) and with de Jesus and Nunes (2024c) who found that most Tetun text on the internet comes from news articles.
@@
Education & research material 55.9% 7.57.57.57.5%
Organizational & formal docs 18.0% 17.3%
News article & press release 10.210.210.210.2% 61.2%
Conversation & correspond. 8.78.78.78.7% 4.84.84.84.8%
Literary fiction 4.34.34.34.3% 0.90.90.90.9%
Religious text 2.52.52.52.5% 8.28.28.28.2%
Other 0.40.40.40.4% 0.30.30.30.3%
Table 4: Provenance distribution in MT and MADLAD.
@@ 4.3.3 Genre
We observe that official (INL) spelling is only loosely followed in Tetun MT inputs, with a median of 71% of words spelled correctly, compared to a median of 82% in MADLAD. This is coherent with previous studies that note the high proportion of Tetun texts that do not follow the official spelling, including in the media Greksakova (2018). We also observe that variance in the rate of correct spelling is higher in MT inputs (standard deviation of 13%) than in MADLAD (standard deviation of 8%) – see Figure 5.
@@
4.4 Effects of domain mismatch on LM modelling
After uncovering a lack of similarity between MT inputs and available corpora in the previous section, we investigate downstream effects on language modelling, by comparing perplexity on MT Tetun inputs with perplexity on monolingual corpora using Goldfish Tetun [^9] https://huggingface.co/goldfish-models/tet_latn_full Chang et al. (2024), a monolingual Tetun language model trained mostly (> 98%) on MADLAD Tetun documents. We use the HuggingFace evaluate [^10] https://github.com/huggingface/evaluate, version 0.4.3 library for all perplexity measurements, and only measure on inputs of 5 words or more.
MT inputs vs monolingual corpora
We compare perplexity on MT Tetun inputs with perplexity on the Labadain Tetun corpus de Jesus and Nunes (2024c), which is the second largest Tetun corpus after MADLAD, and was not used when training the Goldfish model. We also exclude Labadain sentences that are also present in MADLAD. We get an average perplexity of 1,77417741,7741 , 774 on MT Tetun inputs, compared to an average of $153153$ for Labadain, which illustrates that MT inputs may contain vocabulary, structures, or topics that are underrepresented in scraped monolingual corpora, leading to lower language modelling performance and potentially increased translation errors.
Perplexity per topic
Comparing average perplexity per topic, we find that documents in the “Government & socio-economic issues” topic have a much lower perplexity (average of $632632$) than documents from the “Technology”, “Business & work & employment”, or “Daily life & personal experiences” domains (3,43434343,$43$ , 434, 1,87118711,8711 , 871 and 1,86418641,8641 , 864 respectively). This result illustrates that domain imbalance in monolingual corpora has downstream effect on language modelling ability for under-represented domains.
Perplexity per provenance
Comparing average perplexity per provenance, we find that documents that come from “News article & press release” have the lowest perplexity at $874874$, compared to 1,32013201,3201 , 320 for “Education & research material”, 1,66216621,6621 , 662 for “Organizational & formal documents” and 1,86518651,8651 , 865 for “Conversation & correspondence”. Here also, this result shows that document provenances that are under-represented in available corpora impact language modelling performance.
We report all perplexity results in Appendix B.
@@ 5 Recommendations
Anchoring MT evaluation in user demand.
Our results highlight the importance of: (1) broad domain coverage that includes education material, especially about science and health (2) accuracy when translating into, rather than from, the low-resource language, and (3) accuracy when translating very short documents (including individual words). This could start with an evaluation framework that finds a balance between these criteria. In particular, the sharp imbalance between the domain of available Tetun corpora and MT inputs leads us to recommend the systematic evaluation of MT accuracy in a wide variety of domains, and advise against measuring MT accuracy on domains that are over-represented in available corpora (e.g., news or religion), as this would risk giving a poor representation of end-user experience.
Meeting end-users where they are.
We recommend a focus on meeting the needs of students who use MT on their mobile device. Given that students seem to represent a large portion of MT users, and that learning outcomes are an important social good, future study of MT perception and improvement for students in low-resource contexts would be beneficial, both to the research community and end users. Similarly, given the high cost and low reliability of internet connections in Timor-Leste The Asia Foundation (2023), we recommend a heightened focus on MT inference on mobile devices, for example through further research into efficient MT, or through the creation of an open source framework for on-device MT that targets mobile devices. [^11] Similar to the Bergamot / Browser MT project
@@ 6 Conclusion
We present an analysis of MT usage for Tetun, a low-resource language, based on server logs from tetun.org, a widely-used MT service. Our study reveals key insights into user demographics, device preferences, translation directions, and domain coverage. We discover that a large share of users are likely students in Timor-Leste using mobile devices, translating education material from high-resource languages into Tetun. Our analysis also uncovers a significant mismatch between the domains of user-translated text and existing Tetun corpora.
@@
Our recommendations include a stronger emphasis on domains that are not well covered by existing corpora (in particular health and science education material), accuracy for the high-resource $\rightarrow$ low-resource direction, performance on very short texts, and the development of mobile-optimized MT solutions. For human-computer interaction (HCI), we suggest focusing on MT perception and improvement for students in low-resource contexts.
@@
This study provides the first observational analysis of real-world MT usage for a low-resource language. It contributes to a more nuanced understanding of the challenges and opportunities in this field. We hope these insights will guide future research efforts, ultimately leading to more effective and accessible MT solutions for speakers of low-resource languages.
Limitations
System design decisions impacting numbers
Design decisions made by the tetun.org volunteers might have impacted usage numbers. For example, apart from MT, tetun.org includes a multilingual word dictionary (with entry definitions and example sentences), and a Tetun spell checker, which might skew the number of users upward, with potentially some users using the app or website without using the MT functionality. Similarly, the authors chose to only allow translation between Tetun and three higher-resource languages (English, Portuguese, Indonesian), which raises the question of how adding support for more languages would impact usage and balance between languages.
Noise in reported data
The methods used to determine user location may not accurately reflect the actual location and language background of the users. For instance, Timorese users living abroad might have devices registered outside Timor-Leste, and some users may access tetun.org through a VPN. These potential inaccuracies may affect our understanding of user demographics. Additionally, server logs only capture successful translation attempts, introducing potential bias: users may avoid tetun.org for domains or language pairs where it performs poorly, skewing the data toward areas where tetun.org is more effective.
Generalizability
This paper focuses on Tetun, which might limit the applicability of our findings to other languages, given the variability of use cases, user needs and data availability within low-resource communities Liu et al. (2022). While some of our findings are applicable to other low-resource languages, in particular those that are also institutionalized lingua franca, they are less likely to be applicable to much lower-resourced languages, in particular oral and endangered ones Bird and Yibarbuk (2024).
Ethical considerations
Data privacy
While working with real end-user data poses privacy risks, we mitigate them through multiple strategies: (1) Server logs do not include user internet protocol (IP) addresses or personal account information; (2) To protect user confidentiality, we apply Named Entity Recognition using spaCy’s xx_ent_wiki_sm model to systematically remove names and addresses from translation inputs and outputs before analysis; (3) Our analysis exclusively reports aggregated, anonymized results, ensuring no individual user data can be traced or identified.
Ethics application and user consent
We applied for and received a waiver of consent from our institution’s ethics board, as obtaining explicit user consent was not possible given the anonymous nature of the service. This waiver was granted based on several factors: (1) users agree to research use of translation text under the service’s privacy policy, which states: "We reserve the right to use translated text for research and product improvement purposes. All translated text is processed anonymously, ensuring that no personal identifiers are linked to the content"; (2) the service does not collect user accounts or identifying information; and (3) all analysis is conducted on aggregated data. Data management followed strict institutional protocols - all data was transferred directly from the service’s server to our institution’s high-performance computing environment, with no cross-border data flows, and access was restricted to authorized researchers only.
@@ Appendix A Prompt for domain classification
Appendix B Perplexity results
Topic Average perplexity
Science & research 1,20812081,2081 , 208
Healthcare & medicine 1,19811981,1981 , 198
Education 1,37213721,3721 , 372
Business & work & empl. 1,87118711,8711 , 871
Daily life & personal exp. 1,86418641,8641 , 864
Government & socio. issues $632632$
Law & regulations 1,27412741,2741 , 274
Technology 3,43434343,$43$ , 434
Religion & spirituality 1,12711271,1271 , 127
Table 5: Perplexity per topic, as measured on Tetun MT input documents of 5 words or more.
Provenance Average perplexity
Education & research material 1,32113211,3211 , 321
Organizational & formal docs 1,66216621,6621 , 662
News article & press release $874874$
Conversation & correspond. 1,86618661,8661 , 866
Literary fiction 2,11921192,1192 , 119
Religious text 1,47314731,4731 , 473
Other 3,87638763,8763 , 876
Table 6: Perplexity per provenance, as measured on Tetun MT input documents of 5 words or more.